OpenStudy (anonymous):

Which expression is the simplified form of the square root of z to the seventh power? z times the square root of z cubed z squared times the square root of z z cubed times the square root of z z cubed times the square root of z squared

OpenStudy (harsimran_hs4):

square root of z is z^(1/2) then this is raised to 7th power so expression becomes (z^(1/2)) ^ 7 clear ?

OpenStudy (anonymous):

two goes in to 7 3 times with a remainder of 1, so \(x^3\) comes out of the radial and on \(x\) stays in \[\sqrt{x^7}=x^3\sqrt{x}\]

OpenStudy (anonymous):

as is almost always the case (8 times out of 10) the answer is C

OpenStudy (anonymous):

Thank youu!

OpenStudy (anonymous):

yw